# Guido Franchetti1 and Derek Harland2

### v1

@@ 4 Killing spinors and pluricomplex structures
The gauge fixing condition (25) was motivated by a connection with Killing spinors [11]. We will present a precise formulation of this correspondence below, and show that it induces geometric structures previously studied by Nash and Bielawski–Schwachhofer. We then show that these geometric structures are compatible with the symmetric bilinear forms $g_{\pm}$.
@@
To relate this to monopoles, let us suppose that $M$ is a three-dimensional. Then the Clifford algebra admits two inequivalent representations; we choose the one in which the volume form acts as the identity, and this choice determines the spinor bundle $S$. We recall that $S$ is a complex rank 2 bundle admitting a non-degenerate skew-symmetric pairing (,)($\cdot$,$\cdot$)( ⋅ , ⋅ ). Such a pairing induces an isomorphism $S\cong S^{\ast}$, written as $s\mapsto s^{\ast}$, such that the pairing is $(s,t)=s^{\ast}(t)$. In a suitable local frame the map $s\mapsto s^{\ast}$ takes the form
@@
Three-dimensional hyperbolic space $H^{3}$ admits Killing spinors with Killing constant ±$\pm\frac{i}{2}$. The space $K^{\pm}$ of Killing spinors with Killing constant ±$\pm\frac{i}{2}$ is two-dimensional. Since the spinor bundle $S$ is two-dimensional and Killing spinors are determined by their value at a point, Killing spinors trivialise $S$. We thus have the following natural isomorphisms of vector bundles, for $p\in H^{3}$, $\psi^{\pm}\in K^{\pm}$,
@@
More explicitly, this means that given a pair of independent Killing spinors $\psi_{1},\psi_{2}$ in $K^{-}$, any section $(a,\phi)$ of $(\Lambda^{1}\oplus\Lambda^{0})\otimes\mathbb{C}$ can be written uniquely as in the form $a+\phi=\nu_{1}\otimes\psi_{1}^{\ast}+\nu_{2}\otimes\psi_{2}^{\ast}$ for a pair of sections $\nu_{1},\nu_{2}$ of $S$ (and similarly for $K^{+}$).
@@
Let $(A,\Phi)$ be a hyperbolic monopole and let $D^{A}:\Gamma(S\otimes\operatorname{End}(V))\to\Gamma(S\otimes\operatorname{End}(V))$ be the Dirac operator associated to $A$. Under the the isomorphism (79), we have an identity
@@
$-e^{i}\cdot\nabla^{A}_{i}({\phi}+a)=-e^{i}\nabla_{i}^{A}{\phi}-e^{i}\wedge\nabla_{i}^{A}a+e^{i}\lrcorner\nabla_{i}^{A}{a}=-d^{A}{\phi}-d^{A}{a}+*d^{A}*{a}.$ (85)
@@
Proposition (11) gives a relationship between eigenspinors $\nu$ solving
$D^{A}\nu-[\Phi,\nu]=\pm\frac{i}{2}\nu$ (87)
and solutions $(a,\phi)$ of (20), (25) that describe tangent vectors to the moduli space. To make this relationship precise, we need to impose boundary conditions on $\nu$. Just as we did with tangent vectors $(a,\phi)$, we will impose these boundary conditions by lifting to $S^{4}$.
Recall that the spinor bundle of $S^{1}\times H^{3}$ takes the form $S_{+}\oplus S_{-}$. Self-dual two-forms act trivially on $S_{-}$, while anti-self-dual 2-forms act trivially on $S_{+}$. Both $S_{+}$ and $S_{-}$ are isomorphic to the pull-back of the spinor bundle $S$ of $H^{3}$. The Dirac operator associated to the connection $\Phi\,d\theta+A$ splits as the sum of an operator $D^{B}_{H}:\Gamma(S^{1}\times H^{3},S_{-}\otimes\operatorname{End}_{0}(V))\to\Gamma(S^{1}\times H^{3},S_{+}\otimes\operatorname{End}_{0}(V))$ and its adjoint. We focus on $D^{B}_{H}$, which takes the form
@@
extends smoothly to $S^{4}$. Then $E_{x}$ is naturally isomorphic to the kernel of $D^{B}_{S}:\Gamma(S^{4},S_{-}\otimes\operatorname{End}_{0}(V))\to\Gamma(S^{4},S_{+}\otimes\operatorname{End}_{0}(V))$.
By construction, the spinor $\tilde{\nu}$ transforms with weight ±$\pm 1/2$ under the circle action. This fact allows us to determine the dimension of $E_{x}^{\pm}$ using the equivariant index theorem:
@@
To calculate the equivariant Chern character, recall that $V\to S^{2}$ is isomorphic to $L^{+}\oplus L^{-}$ with $L^{-}=(L^{+})^{\ast}$, and so the bundle $\operatorname{End}_{0}(V)\to S^{2}$ is isomorphic to $(L^{+})^{\otimes 2}\oplus\mathbb{C}\oplus(L^{-})^{\otimes 2}$. The degrees of these bundles are $2n,0,-2n$ and their weights under the $U(1)$ action are $-4p,0,4p$. We choose a connection on $L^{+}$ with curvature $-i\omega/2$, where $\omega$ is the area form on $S^{2}$ whose integral is $4\pi$. Then
@@
The normal bundle $NS^{2}$ is topologically trivial, and so $S^{\pm}(NS^{2})$ are also trivial. The action of the circle group on $NS^{2}$ has weights ±$\pm 1$, and the action of its double cover $U(1)$ has weights ±$\pm 2$, so the action of $U(1)$ on the positive and negative spinor bundles has weights ±$\pm 1$. Thus:
@@
The spaces $E_{x}^{\pm}$ are the weight spaces in $\ker D^{B}_{S}$ with weight ±$\pm 1$ under the double cover $U(1)$ of the circle group. Since the connection $B$ is anti-self-dual, a well-known argument shows that the cokernel of $D^{B}_{S}$ is trivial [3]. Therefore the equivariant index (92) is equal to the character of the representation $\ker(D^{B}_{S})$ of $U(1)$. The dimensions of the weight spaces $E^{\pm}_{x}$ are the coefficients $\gamma^{\pm 1}$ in this polynomial, and from the calculation (98) they are both equal to $2n$ as claimed. ∎
Returning to the discussion of boundary conditions, let $\nu\in E_{x}^{-}$ and let $\psi\in K^{-}$. From these we construct a solution $(a_{-},\phi_{-})$ on $S^{1}\times H^{3}$ of (20) and (25) via the isomorphism (77). We wish to show that $b_{-}:=\phi_{-}\,d\theta+a_{-}$ extends over $S^{2}$ to give a smooth section of $\Lambda^{1}S^{4}\otimes\operatorname{End}_{0}(V)$. Using the Clifford isomorphism
@@
induced by the metric $g_{S}$ on $S^{4}$, $b_{-}$ can be thought of as a section of $S_{-}\otimes S_{+}\otimes\operatorname{End}_{0}(V)$, defined away from $S^{2}\subset S^{4}$. The Clifford isomorphisms (77), (99) are related by
$\operatorname{cl}_{S}=\frac{1+|w|^{2}}{2|w|}\operatorname{cl}_{H},$ (100)
in which we have implicitly identified $\Lambda^{1}(S^{4}\setminus S^{2})$ with $\Lambda^{1}(H^{3})\oplus\Lambda^{0}(H^{3})$ and identified $S$ with $S_{+}$ and $S_{-}$. So $\operatorname{cl}_{H}(b_{-})$ is given by
$\left(\frac{1+|w|^{2}}{2|w|}\right)\nu\otimes\psi,$ (101)
@@
$\tilde{\psi}:=\left(\frac{2|w|}{1+|w|^{2}}\right)^{\frac{1}{2}}e^{i\frac{\theta}{2}}\psi.$ (102)
@@
We do so in the upper half-space model of $H^{3}$. Define coordinates $\rho>0,x,y$ by stereographic projection:
@@
We choose a frame for the spinor bundle so that the Clifford isomorphism (77) is given by Pauli matrices:
@@
This ensures that $\mathrm{Vol}\mapsto 1$. Then the Levi-Civita connection on the spinor bundle is
@@
$[\nabla,i\sigma_{1}]=\frac{dx}{\rho}i\sigma_{3},\quad[\nabla,i\sigma_{2}]=\frac{dy}{\rho}i\sigma_{3},\quad[\nabla,i\sigma_{3}]=-\frac{dx}{\rho}i\sigma_{1}-\frac{dy}{\rho}i\sigma_{2}.$ (107)
@@
are Killing spinors with Killing constant $-\frac{\mathrm{i}}{2}$ which form a basis for $K^{-}$.
Now we aim to write $\psi_{\alpha}$ as a spinors on $S^{4}\setminus S^{2}$. It is natural to identify $S^{1}\times H^{3}$ with $\mathbb{R}^{4}\setminus\mathbb{R}^{2}$ using polar coordinates $(\rho\cos\theta,\rho\sin\theta,x,y)$. We choose a frame for the spinor bundle of $S^{4}$ over $\mathbb{R}^{4}$ such that the Clifford isomorphism (99) is given by
@@
in our chosen representations. By comparing this with (100) we can see how our frames for $S$ and $S_{\pm}$ are related. Thus the Killing spinors $\psi_{\alpha}$ written in the frame (105) for $S$ correspond to $\exp(-\frac{i\theta}{2}\sigma_{3})\psi_{\alpha}$ in the frame (110) for $S_{+}$. Then from (102) and (109) we find
@@
These are smooth functions on $\mathbb{R}^{4}$ and they are written in a frame that extends over $\mathbb{R}^{2}\subset\mathbb{R}^{4}$, so we conclude that the spinors $\tilde{\psi}_{\alpha}$ extend smoothly over $\mathbb{R}^{2}\subset\mathbb{R}^{4}$. Repeating the calculation in a different coordinate patch shows that the Killing spinors extend smoothly over $S^{2}\subset S^{4}$.
So if $\nu\in E_{x}^{-}$ and $\psi\in K^{-}$ then $(a_{-},\phi_{-}):=\operatorname{cl}_{H}^{-1}(\nu\otimes\psi^{\ast})$ is a solution of the linearised Bogomolny equations (20) and the gauge fixing condition (25) that extends smoothly over $S^{2}\subset S^{4}$. Similarly, $E^{+}\otimes K^{+}$ gives solutions of the “+” gauge fixing condition. We thus have natural maps to the complexified tangent space:
@@
We now wish to show that these maps are isomorphisms.
We begin by considering surjectivity. By Theorem 4, any tangent vector can be represented by a pair $(a_{-},\phi_{-})$ solving equations (20) and (25). We wish to write this pair in the form
@@
for some eigenspinors $\nu_{i}$ and Killing spinors $\psi_{i}$. We begin by choosing a pair of linearly independent Killing spinors. The Killing spinor equation implies that the function $(\psi_{1},\psi_{2})$ is constant; without loss of generality we can assume that this constant is 1 (it is certainly 1 for the choice made in (109)). Then $\psi_{2}\otimes\psi_{1}^{\ast}-\psi_{1}\otimes\psi_{2}^{\ast}\in S\otimes S^{\ast}$ corresponds to the identity map acting on $S$. So (114) holds with
@@
From equations (91), (100) and (102),
@@
$=\frac{1+|w|^{2}}{w}\epsilon_{\alpha\beta}\operatorname{cl}_{S}(b_{-})\tilde{\psi}_{\beta}.$ (117)
In order to show that $\tilde{\nu}_{\alpha}$ extends smoothly, we need to show that $\operatorname{cl}_{S}(b_{-})\tilde{\psi}_{\beta}=O(w)$, in which the notation $O(w)$ means $w$ times a function that extends smoothly over $\mathbb{R}^{4}\setminus\mathbb{R}^{2}$. Recall from Lemma 10 that $b_{-}^{0,1}=O(w)$. Turning to $b_{-}^{1,0}$, we deduce from $w/(1+|w|^{2})=\rho e^{i\theta}/(1+\rho^{2}+x^{2}+y^{2})$ that $dw=d(\rho e^{i\theta})+O(w)$. So (110) gives
@@
From this and equation (102) we deduce that
@@
So $\operatorname{cl}_{S}(b_{-})\tilde{\psi}_{\beta}$ equals a smooth section times $w$, and hence $\tilde{\nu}_{\alpha}$ given in (117) extends smoothly across $\mathbb{R}^{2}\subset\mathbb{R}^{4}$. Repeating the calculation in a different coordinate patch shows that $\tilde{\nu}_{\alpha}$ extends smoothly over $S^{2}\subset S^{4}$. So $\nu_{\alpha}\in E_{x}^{-}$, and we have shown that (113) is surjective. Since the domain and target both have dimension $4n$, we conclude that
Theorem 13.
The maps given in (113) are isomorphisms.
A similar argument allows us to show that the framed gauge transformation that solves the gauge-fixing condition in Theorem 4 is unique. Suppose to the contrary: then there is a non-zero framed gauge transformation $\chi$ such that $(a_{-},\phi_{-})=(d^{A}\chi,0)$ solves (20) and (25). This can be written in the form (114) with eigenspinors $\nu_{\alpha}\in E_{x}$ given in (115). Since the map (113) is injective and $(d^{A}\chi,0)$ represents the zero-vector in the quotient (22), it must be that $\nu_{\alpha}=0$. Then, from (114), $d^{A}\chi=0$. Since monopoles are irreducible this in turn implies that $\chi=0$, which contradicts our assumption.
[…]

### v2

@@ 4 Killing spinors and pluricomplex structures
The gauge fixing condition (49) was motivated by a connection with Killing spinors [12]. We will present a precise formulation of this correspondence below, and show that it induces various geometric structures on the framed moduli space. We then show that these geometric structures are compatible with the symmetric bilinear forms $g_{\pm}$.
4.1 Killing spinors and Dirac operators
@@
To relate this to monopoles, let us suppose that $M$ is three-dimensional. Then the Clifford algebra admits two inequivalent irreducible representations; we choose the one in which the volume form acts as the identity, and this choice determines the spinor bundle $S$. We recall that $S$ is a complex rank 2 bundle admitting a non-degenerate skew-symmetric pairing $(\cdot,\cdot)$. Such a pairing induces an isomorphism $S\cong S^{\ast}$, written as $s\mapsto s^{\ast}$, such that the pairing is $(s,t)=s^{\ast}(t)$. In a suitable local frame the map $s\mapsto s^{\ast}$ takes the form
@@
Three-dimensional hyperbolic space $H^{3}$ admits Killing spinors with Killing constant $\pm\frac{i}{2}$. The space $K^{\pm}$ of Killing spinors with Killing constant $\pm\frac{i}{2}$ is two-dimensional [15]. Since the spinor bundle $S$ is two-dimensional and Killing spinors are determined by their value at a point, Killing spinors trivialise $S$. We thus have the following natural isomorphisms of vector bundles, for $p\in H^{3}$, $\psi^{\pm}\in K^{\pm}$,
@@
More explicitly, this means that given a pair of independent Killing spinors $\psi_{1},\psi_{2}$ in $K^{-}$, any section $(a,\phi)$ of $(\Lambda^{1}\oplus\Lambda^{0})\otimes\mathbb{C}$ can be written in the form $(a,\phi)=\operatorname{cl}_{H}^{-1}(\nu_{1}\otimes\psi_{1}^{\ast}+\nu_{2}\otimes\psi_{2}^{\ast})$ for a unique pair of sections $\nu_{1},\nu_{2}$ of $S$ (and similarly for $K^{+}$).
@@
Let $(A,\Phi)$ be a hyperbolic monopole and let $D^{A}:\Gamma(S\otimes\operatorname{End}(V))\to\Gamma(S\otimes\operatorname{End}(V))$ be the Dirac operator associated to $A$. Under the isomorphism (112), we have the identity
@@
$e^{i}\cdot\nabla^{A}_{i}({\phi}+a)=-e^{i}\nabla_{i}^{A}{\phi}-e^{i}\wedge\nabla_{i}^{A}a+e^{i}\lrcorner\nabla_{i}^{A}{a}=-d^{A}{\phi}-d^{A}{a}+*d^{A}*{a}.$ (118)
@@
4.2 Eigenspinors and tangent vectors
Proposition 24 shows that tangent vectors $(a,\phi)$ to the moduli space of hyperbolic monopoles solving (43), (49) give rise to eigenspinors $\nu$ solving
$D^{A}\nu-[\Phi,\nu]=\pm\frac{i}{2}\nu.$ (120)
We would like to establish a precise relationship between tangent vectors and eigenspinors. To do this we must impose boundary conditions on the solutions $\nu$ of (120). We will do so by lifting the spinors to $S^{4}$.
Recall that the spinor bundle of $S^{1}\times H^{3}$ takes the form $S_{+}\oplus S_{-}$, in which $S_{\pm}$ are eigenspaces of the action of the volume form by Clifford multiplication with eigenvalues $\mp 1$. This means that self-dual two-forms act by zero on $S_{-}$, while anti-self-dual 2-forms act by zero on $S_{+}$. Both $S_{+}$ and $S_{-}$ are isomorphic to the pull-back of the spinor bundle $S$ of $H^{3}$. The Dirac operator associated to the connection $\Phi\,d\theta+A$ splits as the sum of an operator $D^{B}_{H}:\Gamma(S^{1}\times H^{3},S_{-}\otimes\operatorname{End}_{0}(V))\to\Gamma(S^{1}\times H^{3},S_{+}\otimes\operatorname{End}_{0}(V))$ and its adjoint. We focus on $D^{B}_{H}$, which takes the form
@@ 4.2 Eigenspinors and tangent vectors
extends smoothly to $S^{4}$. Then $E_{x}^{\pm}$ is naturally isomorphic to the kernel of $D^{B}_{S}:\Gamma(S^{4},S_{-}\otimes\operatorname{End}_{0}(V))\to\Gamma(S^{4},S_{+}\otimes\operatorname{End}_{0}(V))$.
By construction, the spinor $\tilde{\nu}$ transforms with weight $\pm 1/2$ under the circle action. This fact allows us to determine the dimension of $E_{x}^{\pm}$ using the equivariant index theorem:
@@
To calculate the equivariant Chern character, recall that $V\to S^{2}$ is isomorphic to $L^{+}\oplus L^{-}$ with $L^{-}=(L^{+})^{\ast}$, and so the bundle $\operatorname{End}_{0}(V)\to S^{2}$ is isomorphic to $(L^{+})^{\otimes 2}\oplus\mathbb{C}\oplus(L^{-})^{\otimes 2}$. The degrees of these bundles are 2n,0,2n2n,0,-2n and their weights under the $U(1)$ action are 4p,0,4p-4p,0,4p. We choose a connection on $L^{+}$ with curvature $i\omega/2$, where $\omega$ is the area form on $S^{2}$ whose integral is $4\pi$. Then
@@
The normal bundle $NS^{2}$ is topologically trivial, and so $S^{\pm}(NS^{2})$ are also trivial. The action of the circle group on $NS^{2}$ has weights $\pm 1$, and the action of its double cover $U(1)$ has weights $\pm 2$, so the action of $U(1)$ on the positive and negative spinor bundles has weights $\pm 1$. Thus:
@@
The spaces $E_{x}^{\pm}$ are the weight spaces in $\ker D^{B}_{S}$ with weight $\pm 1$ under the double cover $U(1)$ of the circle group. Since the connection $B$ is anti-self-dual, a well-known argument shows that the cokernel of $D^{B}_{S}$ is trivial [3]. Therefore the equivariant index (125) is equal to the character of the representation $\ker(D^{B}_{S})$ of $U(1)$. The dimensions of the weight spaces $E^{\pm}_{x}$ are the coefficients $\gamma^{\pm 1}$ in this polynomial, and from the calculation (131) they are both equal to $2n$ as claimed. ∎
We now aim to relate $E_{x}^{\pm}$ to the tangent space of the framed monopole moduli space. Let $\nu\in E_{x}^{\pm}$ and let $\psi\in K^{\pm}$. Then $\nu\otimes\psi$ gives a section $(a_{\pm},\phi_{\pm})$ of $(\Lambda^{1}H^{3}\oplus\Lambda^{0}H^{3})\otimes\operatorname{End}_{0}(V)$ via the isomorphism (110). Proposition 24 shows that $(a_{\pm},\phi_{\pm})$ solves the linearised Bogomolny equation (43) and the gauge-fixing condition (49). To show that this is a tangent vector to the moduli space we must show the following:
Proposition 26.
If $\nu\in E_{x}^{\pm}$, $\psi\in K^{\pm}$ and $(a_{\pm},\phi_{\pm}):=\operatorname{cl}_{H}^{-1}(\nu\otimes\psi)$, then $b_{\pm}:=\phi_{\pm}\,d\theta+a_{\pm}$ extends over $S^{2}$ to give a smooth section of $\Lambda^{1}S^{4}\otimes\operatorname{End}_{0}(V)$.
Proof.
For notational simplicity we prove the proposition only for $b_{-}$. The proof for $b_{+}$ is similar. Using the Clifford isomorphism
@@
induced by the metric $g_{S}$ on $S^{4}$, $b_{-}$ can be identified with the section $\operatorname{cl}_{S}(b_{-})$ of $S_{-}\otimes S_{+}\otimes\operatorname{End}_{0}(V)$, defined away from $S^{2}\subset S^{4}$. The Clifford isomorphisms (110), (132) are related by
$\operatorname{cl}_{S}=\left(\frac{1+|w|^{2}}{2|w|}\right)\operatorname{cl}_{H},$ (133)
in which we have implicitly identified $\Lambda^{1}(S^{4}\setminus S^{2})$ with $\Lambda^{1}(H^{3})\oplus\Lambda^{0}(H^{3})$ and identified $S$ with $S_{+}$ and $S_{-}$. So $\operatorname{cl}_{S}(b_{-})$ is given by
$\operatorname{cl}_{S}(b_{-})=\left(\frac{1+|w|^{2}}{2|w|}\right)\operatorname{cl}_{H}(a_{-},\phi_{-})=\left(\frac{1+|w|^{2}}{2|w|}\right)\nu\otimes\psi,$ (134)
@@
$\tilde{\psi}:=\left(\frac{2|w|}{1+|w|^{2}}\right)^{\frac{1}{2}}e^{\frac{i\theta}{2}}\psi.$ (135)
@@
We do so in the upper half-space model of $H^{3}$. Recall that $S^{4}$ is the set of unit vectors $y\in\mathbb{R}^{5}$. We define coordinates $\theta,\rho>0,x,y$ on $S^{4}\setminus S^{2}\simeq S^{1}\times H^{3}$ as follows:
@@
We choose a frame for the spinor bundle $S\to H^{3}$ so that the Clifford isomorphism (110) is given by Pauli matrices:
@@
This ensures that $\operatorname{cl}_{H}(\mathrm{Vol})$ is the identity matrix $I_{2}$. Then the Levi-Civita connection on the spinor bundle is
@@
$\nabla,i\sigma_{1}]=\frac{dx}{\rho}i\sigma_{3},\quad[\nabla,i\sigma_{2}]=\frac{dy}{\rho}i\sigma_{3},\quad[\nabla,i\sigma_{3}]=-\frac{dx}{\rho}i\sigma_{1}-\frac{dy}{\rho}i\sigma_{2}.$ (140)
@@
are Killing spinors with Killing constant $\frac{\mathrm{i}}{2}$ which form a basis for $K^{-}$.
Now we aim to write $\psi_{\alpha}$ as a spinors on $S^{4}\setminus S^{2}$. We identify $S^{1}\times H^{3}\simeq S^{4}\setminus S^{2}$ with $\mathbb{R}^{4}\setminus\mathbb{R}^{2}$. The map $S^{1}\times H^{3}\to\mathbb{R}^{4}\setminus\mathbb{R}^{2}$ is given by $(\theta,\rho,x,y)\mapsto(\rho\cos\theta,\rho\sin\theta,x,y)$. The map $S^{4}\setminus S^{2}\to\mathbb{R}^{4}\setminus\mathbb{R}^{2}$ is stereographic projection. The metric $g_{S}$ is then $4/(1+\rho^{2}+x^{2}+y^{2})^{2}$ times the euclidean metric. We can choose frames for $S_{+},S_{-}\to S^{4}$ on the whole of $\mathbb{R}^{4}$ such that the Clifford isomorphism (132) is given by
@@
[…]
@@
These are smooth functions on $\mathbb{R}^{4}$ and they are written in a frame that extends over $\mathbb{R}^{2}\subset\mathbb{R}^{4}$, so we conclude that the spinors $\tilde{\psi}_{\alpha}$ extend smoothly over $\mathbb{R}^{2}\subset\mathbb{R}^{4}$. Repeating the calculation in a different coordinate patch shows that $\tilde{\psi_{\alpha}}$ extend smoothly over $S^{2}\subset S^{4}$. ∎
Proposition 26 shows that there are natural maps
@@
We now wish to show that:
Theorem 27.
The maps given in (146) are isomorphisms.
Proof.
We begin by considering surjectivity. Let $Z_{x}^{-}$ denote the space of solutions $(a_{-},\phi_{-})$ of (43), (49) such that $b_{-}=\phi_{-}\,d\theta+a_{-}$ extends smoothly to $S^{4}$. Then (146) is a composition of maps
$E_{x}^{-}\otimes K^{-}\to Z_{x}^{-}\to T_{x}^{\mathbb{C}}M^{f}_{n,p}.$ (147)
Theorem 13 shows that the map $Z_{x}^{-}\to T_{x}^{\mathbb{C}}M^{f}_{n,p}$ is surjective. We now show that the map $E_{x}^{-}\otimes K^{-}\to Z_{x}^{-}$ is surjective.
Let $(a_{-},\phi_{-})\in Z_{x}^{-}$. We wish to write this in the form
@@
for some eigenspinors $\nu_{i}\in E_{x}^{-}$ and Killing spinors $\psi_{i}\in K^{-}$. We begin by choosing a pair of linearly independent Killing spinors. The Killing spinor equation implies that the function $(\psi_{1},\psi_{2})$ is constant; without loss of generality we can assume that this constant is 1 (it is certainly 1 for the choice made in (142)). Then $\psi_{2}\otimes\psi_{1}^{\ast}-\psi_{1}\otimes\psi_{2}^{\ast}\in\Gamma(S\otimes S^{\ast})$ corresponds to the identity map acting on $S$. So (148) holds with
@@
Let $\tilde{\psi}_{\alpha}$ be constructed from $\psi_{\alpha}$ as in (135). From equations (124), (133) and (135),
@@
$=\left(\frac{1+|w|^{2}}{w}\right)\epsilon_{\alpha\beta}\operatorname{cl}_{S}(b_{-})\tilde{\psi}_{\beta}.$ (151)
In order to show that $\tilde{\nu}_{\alpha}$ extends smoothly, we need to show that $\operatorname{cl}_{S}(b_{-})\tilde{\psi}_{\beta}=O(w)$, in which the notation $O(w)$ means $w$ times a function that extends smoothly over $S^{2}\subset S^{4}$. We do this calculation using the identification $S^{1}\times H^{3}\simeq S^{4}\setminus S^{2}$ with $\mathbb{R}^{4}\setminus\mathbb{R}^{2}$ that was introduced in the proof of Proposition 26.
Recall from Lemma 23 that $b_{-}^{0,1}=O(w)$. Turning to $b_{-}^{1,0}$, we deduce from $w/(1+|w|^{2})=\rho e^{i\theta}/(1+\rho^{2}+x^{2}+y^{2})$ that $dw=d(\rho e^{i\theta})+O(w)$. So (143) gives
@@
From this and equations (135), (142) we deduce that
@@
So $\operatorname{cl}_{S}(b_{-})\tilde{\psi}_{\beta}$ equals a smooth section times $w$, hence $\tilde{\nu}_{\alpha}$ given in (151) extends smoothly across $\mathbb{R}^{2}\subset\mathbb{R}^{4}$. Repeating the calculation in a different coordinate patch shows that $\tilde{\nu}_{\alpha}$ extends smoothly over $S^{2}\subset S^{4}$. So $\nu_{\alpha}\in E_{x}^{-}$, and we have shown that (146) is a surjective linear map. Since the domain and target both have dimension $4n$, it is an isomorphism. ∎
We are now in a position to prove the uniqueness part of Theorem 13:
Proposition 28.
Let $(A,\Phi)$ and $(a,\phi)$ be as in Theorem 12. Let $\chi^{(1)},\chi^{(2)}$ be two infinitesimal unframed $\mathfrak{sl}(2,\mathbb{C})$ gauge transformations such that $(a_{-}^{(p)},\phi_{-}^{(p)})=(a+d^{A}\chi^{(p)},\phi+[\Phi,\chi^{(p)}])$ solve the gauge fixing condition (49). Then $\chi^{(1)}=\chi^{(2)}$.
Proof.
The statement of the proposition is equivalent to injectivity of the map $Z_{x}\to T_{x}^{\mathbb{C}}M_{n,p}^{f}$ in (147). The proof of Theorem 27 showed that the map $E_{x}^{-}\otimes K^{-}\to Z_{x}^{-}$ is surjective. Since $\dim(E_{x}^{-}\otimes K^{-})=4n$, this means that $\dim(Z_{x}^{-})\leq 4n$. Corollary 11 shows that $\dim(T_{x}^{\mathbb{C}}M^{f}_{n,p})=4n$, so
$\dim(Z_{x}^{-})\leq\dim(T_{x}^{\mathbb{C}}M^{f}_{n,p}).$ (154)
The existence part of Theorem 13 (which was proved in section 3) says that the map $Z_{x}^{-}\to T_{x}^{\mathbb{C}}M_{n,p}^{f}$ is surjective. Since $\dim(Z_{x}^{-})\leq\dim(T_{x}^{\mathbb{C}}M^{f}_{n,p})$, this map must also be injective, by the rank-nullity theorem. ∎
4.3 Symplectic and pluricomplex structures
Now we discuss some some geometric structures that can be derived from the tensor decomposition (146) of the tangent space. Fix a non-zero $\psi\in K^{+}$, and let $\langle\psi\rangle\in\mathbb{P}K^{+}$ be the line through $\psi$. We will first show that
Proposition 29.
There is a unique almost complex structure $J_{\langle\psi\rangle}$ on $TM^{f}_{n,p}$ whose holomorphic tangent bundle is
$T^{1,0}_{\langle\psi\rangle}M^{f}_{n,p}=E^{+}\otimes\langle\psi\rangle.$ (155)
To prove this we will need to make use of the quaternionic structure on the spinor bundle. The spinor bundle $S\to H^{3}$ admits a hermitian metric $\langle\cdot,\cdot\rangle$ and a quaternionic structure, i.e. an antilinear map $S\to S$, $s\mapsto\bar{s}$ that squares to 1-1. These are related to the symplectic pairing by the identities $\langle s,t\rangle=(\bar{s},t)=\bar{s}^{\ast}(t)$. In a suitable local frame, they can be written